OpenStudy (anonymous):

Solve the system by substitution 2x + y = -11 3x -4y = 11 Needing help with substitution..

OpenStudy (anonymous):

get one of the variables to the same coefficient

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I mis - read the question !!

OpenStudy (anonymous):

2x + y = -11 move 2x to the other side

OpenStudy (anonymous):

so y= -11 - 2x?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

now, you can substitute y in the 2nd eq 3x -4y = 11 3x - 4( -11 -2x ) = 11

OpenStudy (anonymous):

So pretty much I just have to solve the first equation and plug it in to the second one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright. Sounds easy enough.. Thank you(: